$SPAD/src/input e02bbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ncap7:=11
lamda:Matrix SF:=
 [[1.00 ,1.00 ,1.00 ,1.00 ,3.00 ,6.00 ,_
   8.00 ,9.00 ,9.00 ,9.00 ,9.00 ]]
c:Matrix SF:=
 [[1.00 ,2.00 ,4.00 ,7.00 ,6.00 ,4.00 ,_
   3.00 ,0.00 ,0.00 ,0.00 ,0.00 ]]
x:=2.0
result:=e02bbf(ncap7,lamda,c,x,-1)

———-

References^{[1] nothing}

4